# LEAP-VO: Long-term Effective Any Point Tracking for Visual Odometry

### v2

@@ 1 Introduction
Visual odometry (VO) calculates the camera’s position and movement from an image sequence. It is widely used in various fields including robotics [46], mixed reality [23], and autonomous driving [15]. The performance of VO highly depends on the accuracy of its point tracking front-end, which recovers relative camera motion between consecutive frames using projective geometry. It further requires the point correspondences to be static, as dynamic points can lead to inaccurate motion estimation and pose recovery.
@@
Recently, Persistent Independent Particles (PIPs) [16] proposes a novel pipeline that brings the idea of iterative refinement into the long-term point tracking task, with the local correlation evidence to maintain feasible computation. While it is proficient in leveraging the temporal context present in the image sequence, it ignores the inter-track information in the image sequence and processes each query point independently. Consequently, PIPs struggles to capture the global motion patterns from the image sequence. Additionally, it lacks a reliability metric for its trajectory estimates, hindering its application in quality-sensitive tasks like video segmentation and visual odometry. To mitigate the aforementioned issues, we propose a pipeline that explores dynamic track estimation and temporal probabilistic formulation for long-term point tracking.
@@ 2 Related Work
In the realm of long-term point tracking or Tracking Any Point (TAP), various methods have been introduced recently. PIPs [16] reframes pixel tracking as a long-term motion estimation with local information and improves tracking accuracy by associating each pixel with a trajectory that spans multiple image frames. The subsequent works enhance performance by integrating spatial context features [2], tracking multiple queries concurrently [18], and extending temporal fields and templates [49]. TAP-Net [9] adopts a different method by locating candidate correspondences through a global cost volume, improved by TAPIR’s trajectory refinement [10] and BootsTAP’s semi-supervised training [11]. Other approaches explore self-supervision targets to learn visual correspondence via powerful feature representation [17, 5, 36]. Recently, OmniMotion [37] provides a pure test-time optimization method that infers the point trajectory and occlusion directly from videos with optical flow predictions. However, its slow inference speed makes it unsuitable for real-time applications such as visual odometry.
@@
Visual Odometry (VO) can be categorized into indirect (feature-based) and direct methods. Indirect methods, such as MonoSLAM [8] and ORB-SLAM [25], mostly rely on feature extraction and matching to estimate the camera pose. Direct methods, conversely, skip feature extraction and matching and utilize intensity to optimize camera pose and 3D scene point positions based on photometric errors [12, 13, 38], showing better performance on low-texture regions where keypoint tracking can easily fail. However, they are more vulnerable to illumination changes and issues with rolling shutter cameras. Advancements in deep learning have spawned end-to-end monocular VO methods in supervised [39, 3, 30, 43, 40, 33, 28, 44] and unsupervised [51, 45, 27, 21] settings. DeepVO [39] utilizes Recurrent Neural Networks to model sequential data, while SfMLearner [51] develops an unsupervised framework to learn the depth and camera motion from unlabeled monocular videos. TartanVO [40] proposes a VO model that generalizes to multiple datasets and real-world scenarios. DROID-SLAM [32] follows learning-based optimization from RAFT [31] and proposes an end-to-end system, integrating flow, confidence, and geometric optimization via iterative GRU updates. The followed-up work DPVO [33] further cuts computational complexity by replacing dense feature tracking with sparse patch tracking.
@@
In addition to enhancing accuracy, numerous studies have focused on the detection of dynamic objects within videos. Methods based on segmentation, such as DynaSLAM [1] and RCVD [20], employ semantic cues from pre-trained models to exclude dynamic regions. However, these segmentation-based methods are constrained to predefined object categories and struggle to capture real-world motion (for example, differentiating between a stationary and a moving car). Alternatively, DytanVO [28] and ParticleSfM [48] adopt a trajectory-based motion detection approach. This method addresses dynamic scenes using estimated point track information, thereby generating reliable camera trajectories in complex motion scenarios. Nonetheless, they depend on dense optical flow estimation to capture global motion patterns. Furthermore, dynamic tracks can be detected using additional depth information for a 3D consistency check [7, 47], or by optimization based on photo consistency [35, 41, 42].
@@ 3.1 Preliminary: Tracking Any Point
Tracking arbitrary pixels in an image sequence is traditionally akin to estimating optical flow, where each pixel is mapped to a corresponding location in a subsequent frame. However, two-view approaches often overlook the rich temporal context within the image sequence, leading to suboptimal outcomes. Recent advancements address this by revisiting the particle video concept and improving long-term point tracking through learning-based techniques [16, 10, 18]. These TAP methods can track points over multiple frames and detect occlusions, significantly boosting tracking robustness under challenging scenarios.
@@
PIPs [16], a notable learning-based TAP method, offers a promising solution featuring local cost volumes and iterative refinement. Initially, each image $\mathbf{I}_{s}\in\mathbf{I}$ is processed through a CNN feature extractor $\mathcal{F}$ to extract the feature map $\mathbf{Y_{s}}=\mathcal{F}(\mathbf{I}_{s})$. Point features $\mathbf{f}_{q}$ are computed by bilinear sampling at the query positions $\mathbf{x_{s}}$ on $\mathbf{Y}_{s}$. These point features from each image are then concatenated to form a point feature tensor $\mathbf{F}=[\mathbf{f}_{1},\ldots,\mathbf{f}_{S}]$. PIPs approaches the TAP problem by iteratively updating the state variables $(\mathbf{X},\mathbf{F})$, with the initial states obtained by duplicating the query positions and features. During the $k$-th iterative refinement, PIPs computes a multi-scale local cost volume $\mathbf{C}[\mathbf{X}^{k}]$ centered around the current estimated point position $\mathbf{X}^{k}$. Utilizing this local evidence, PIPs predicts updates for the state variables via
@@
where $\text{pos}(\cdot)$ refers to the positional embedding. The final point feature $\mathbf{F}^{K}$ is used to predict point visibility with a simple linear projection layer $\mathcal{G}_{v}$, yielding $\mathbf{V}=\mathcal{G}_{v}(\mathbf{F}^{K})$. Recent advances extend PIPs by replacing the MLP-based refiner with an attention-based refiner and adding spatial feature aggregation to track multiple queries together [18].
@@ 3.2 Anchor-based Dynamic Track Estimation
We propose an innovative method for dynamic track estimation that leverages visual, temporal, and inter-track information. While common techniques like monocular segmentation [20] excel at identifying moving objects using visual cues but can erroneously label static objects, like parked cars, as dynamic. Trajectory-based methods [48], utilizing optical flows from two or multiple frames, can differentiate static and moving objects apart but tend to ignore visual information and can be inefficient due to dense trajectories computation. Merging these two distinct methods is nontrivial; however, the formulation of the TAP pipeline, enhanced with inter-track attention, offers a viable solution.
@@ 3.3 Temporal Probability Modeling
Supervised learning applied to large-scale datasets has yielded notable advancements in learning-based point tracking methods. Nonetheless, trajectory estimation errors remain a concern, especially under challenging conditions like occlusions, motion blur, or textureless regions. Recognizing the varying reliability of trajectories, integrating an uncertainty estimation mechanism into the correspondence estimation process is beneficial [6]. This is particularly crucial for quality-sensitive tasks such as visual odometry.
@@
Kernel-based Estimation. Rather than just predicting the positions of points within a trajectory, our model focuses on recovering the distribution of point trajectories by estimating its parameters. The location parameter predictions can be associated with the previous point trajectory as $[\bm{\mu}_{a};\bm{\mu}_{b}]=\mathbf{X}$. However, directly deriving a symmetric and positive definite scale matrix from model outputs is challenging. To address this, we employ a kernel-based approach with the linear kernel $K(\mathbf{x},\mathbf{y})=\mathbf{x}^{T}\mathbf{y}$. During each iteration, two scale matrices $(\mathbf{\Sigma}_{a},\mathbf{\Sigma}_{b})$ are constructed by first applying two linear projection layers to the point features, represented as $\mathbf{F}_{a}^{k}=\mathcal{G}_{a}(\mathbf{F}^{k})$ and $\mathbf{F}_{b}^{k}=\mathcal{G}_{b}(\mathbf{F}^{k})$. We then compute the scale matrices as $\mathbf{\Sigma}_{a}=K(\mathbf{F}_{a},\mathbf{F}_{a})+\sigma\mathbf{I}$, and $\mathbf{\Sigma}_{b}=K(\mathbf{F}_{b},\mathbf{F}_{b})+\sigma\mathbf{I}$, where $\sigma$ is a small positive value and $\mathbf{I}$ is the identity matrix. The model parameters are refined through Maximum Likelihood Estimation (MLE) using a negative log-likelihood (NLL) loss function.
@@ 3.4 LEAP-VO
Keypoint Extraction and Tracking. In our VO system, the LEAP model serves as the front-end. When a new image frame $\mathbf{I}_{t}$ is captured at time $t$, we sample $N$ new keypoints from this image to initiate point tracking. We consistently employ the distributed image gradient-based sampling method (refer to Section 3.2) for keypoint selection due to its favorable properties in feature distinctiveness and distribution. This approach also obviates the need for additional anchors during tracking. The LEAP model then tracks these keypoints across the last $S_{LP}$ frames, performing bidirectional tracking (both forward and backward in time). Notably, $S_{LP}$ can exceed the model window $S$, thus enabling more expansive tracking capabilities through sequential prediction [16, 18].
Track Filtering. Theoretically, precise camera pose estimation between two images can be attained with just a few accurate correspondences under certain conditions. However, the quality of correspondences obtained from feature matching or optical flow can vary. Traditional VO pipelines often use RANSAC [14] to filter out incorrect matches, but their efficiency decreases with more keypoints and views. Our approach, in contrast, bypasses the sampling-based method and uses trajectory quality assessments from LEAP for more efficient and informative track filtering.
@@
where $\|\cdot\|_{\rho}$ is the distance metric, and the weight $w_{i\rightarrow j,n}$ is derived from the track filtering results. We use the Gauss-Newton method [34] to optimize Sec. 3.4 for $K_{BA}$ iterations. At each iteration, we compute the camera poses update $\Delta\xi^{(k)}\in\mathfrak{se}(3)~{}\textrm{(lie-algebra corresponding to}~{}\mathbf{T})$, and the depth update $\Delta\mathbf{D}^{(k)}$ for each point. The optimization can be solved efficiently with the Schur decomposition. We use the robust Huber loss function for the distance metric.
@@ 4.1 Datasets and Metrics
Replica. Replica [29] provides synthetic indoor environments for simulating the image-capturing process with a moving camera. We use the camera trajectories provided by [50] for the localization task, resulting in 16 camera trajectories derived from 8 static scenes. Each sequence comprises 900 frames with rendered RGB images. We use “Sequence 1" for the visual odometry evaluation.
MPI Sintel. The MPI Sintel [4] dataset contains dynamic image sequences from open-source 3D animated short films. These sequences consist of large and complex object motion, along with other imaging effects such as motion blur and defocus blur. Following prior works [20, 48], we assess dynamic VO performance on the MPI Sintel dataset, with each sequence containing 20 to 50 image frames.
TartanAir Shibuya. The TartanAir Shibuya dataset [26] features dynamic scenes from two scenarios: “Standing Humans" and “Road Crossing". The majority of the humans remain stationary in "Standing Humans" sequences, while "Road Crossing" sequences provide a more challenging setting where multiple humans are moving in different directions. Each sequence in the dataset includes 100 frames with more than 30 tracked moving humans.
Metrics. For visual odometry, we compare the Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot) following  [48, 47, 33]. ATE measures how much the estimated trajectory deviated from the ground truth trajectory, which is calculated as the root mean square error (RMSE) overall all pose transformations after scaling and alignment. “RPE trans” computes translation errors made over a certain distance (meter), and “RPE rot” computes the estimated rotation over a certain distance (degree). Both “RPE trans” and “RPE rot” are calculated on all poses and then averaged.
@@
ORB-SLAM2 [24] 0.086 0.030 0.650
DynaSLAM [1] 0.039 0.017 0.366
DROID-SLAM [32] 0.267 0.036 2.631
TartanVO [40] 0.406 0.036 2.063
DytanVO [28] 0.289 0.035 2.146
DPVO [33] 0.257 0.036 2.635
@@
Table 1: Camera tracking results on Replica [29]. The statistics demonstrate that our method achieves better results than other competitive baselines.
@@
ORB-SLAM2 [24] X X X
DynaSLAM [1] X X X
DROID-SLAM [32] 0.175 0.084 1.912
TartanVO [40] 0.238 0.093 1.305
DytanVO [28] 0.131 0.097 1.538
DPVO [33] 0.076 0.078 1.722
@@
Table 2: Camera tracking results on MPI Sintel [4]. The statistics show that our method outperforms other competitive baselines.‘X’ denotes the tracking failure case.
@@
ORB-SLAM2 [24] w/ mask 0.0788 0.0060 0.0657 0.0196 0.0148 1.0984 0.8476 0.3044
DynaSLAM [1] X 0.8836 0.3907 0.4196 0.4925 0.6446 0.6539 (0.5808)
AirDOS [26] w/ mask 0.0606 0.0193 0.0951 0.0331 0.0206 0.2230 0.5625 0.1449
DROID-SLAM [32] 0.0051 0.0073 0.0103 0.0120 0.2278 0.0253 0.5788 0.1238
TartanVO [40] 0.0600 0.1605 0.2762 0.1814 0.2174 0.3228 0.5009 0.2456
DytanVO [28] 0.0327 0.1017 0.0608 0.0516 0.0755 0.0365 0.0660 0.0607
DPVO [39] 0.0539 0.1603 0.1788 0.2858 0.1196 0.1119 0.1498 0.1514
@@
Table 3: ATE (m) results on TartanAir-Shibuya Sequences [26]. ()($\cdot$)( ⋅ ) denotes averaging on valid scenes only. We can observe that our method shows better results for ATE on most sequences, outperforming all SLAM and VO baselines on average ATE.
@@
Figure 5: Qualitative results of camera trajectory estimation on TartanAir-Shibuya [26]. The visualizations show that our method provides more robust and accurate camera pose trajectories, especially in hard cases (RoadCrossing 06 and RoadCrossing 07).

### v3

@@ 1 Introduction
Visual odometry (VO) calculates the camera’s position and movement from an image sequence. It is widely used in various fields including robotics [yousif2015overview], mixed reality [mossel2016streaming], and autonomous driving [geiger2013vision]. The performance of VO highly depends on the accuracy of its point tracking front-end, which recovers relative camera motion between consecutive frames using projective geometry. It further requires the point correspondences to be static, as dynamic points can lead to inaccurate motion estimation and pose recovery.
@@
Recently, Persistent Independent Particles (PIPs) [harley2022particle] proposes a novel pipeline that brings the idea of iterative refinement into the long-term point tracking task, with the local correlation evidence to maintain feasible computation. While it is proficient in leveraging the temporal context present in the image sequence, it ignores the inter-track information in the image sequence and processes each query point independently. Consequently, PIPs struggles to capture the global motion patterns from the image sequence. Additionally, it lacks a reliability metric for its trajectory estimates, hindering its application in quality-sensitive tasks like video segmentation and visual odometry. To mitigate the aforementioned issues, we propose a pipeline that explores dynamic track estimation and temporal probabilistic formulation for long-term point tracking.
@@ 2 Related Work
In the realm of long-term point tracking or Tracking Any Point (TAP), various methods have been introduced recently. PIPs [harley2022particle] reframes pixel tracking as a long-term motion estimation with local information and improves tracking accuracy by associating each pixel with a trajectory that spans multiple image frames. The subsequent works enhance performance by integrating spatial context features [bian2023contexttap], tracking multiple queries concurrently [karaev2023cotracker], and extending temporal fields and templates [zheng2023pointodyssey]. TAP-Net [doersch2022tap] adopts a different method by locating candidate correspondences through a global cost volume, improved by TAPIR’s trajectory refinement [doersch2023tapir] and BootsTAP’s semi-supervised training [doersch2024bootstap]. Other approaches explore self-supervision targets to learn visual correspondence via powerful feature representation [jabri2020space, caron2021emerging, tumanyan2024dino]. Recently, OmniMotion [wang2023omnimotion] provides a pure test-time optimization method that infers the point trajectory and occlusion directly from videos with optical flow predictions. However, its slow inference speed makes it unsuitable for real-time applications such as visual odometry.
@@
Visual Odometry (VO) can be categorized into indirect (feature-based) and direct methods. Indirect methods, such as MonoSLAM [davison2007monoslam] and ORB-SLAM [mur2015orb], mostly rely on feature extraction and matching to estimate the camera pose. Direct methods, conversely, skip feature extraction and matching and utilize intensity to optimize camera pose and 3D scene point positions based on photometric errors [engel2014lsd, engel2017direct, wang2017stereo], showing better performance on low-texture regions where keypoint tracking can easily fail. However, they are more vulnerable to illumination changes and issues with rolling shutter cameras. Advancements in deep learning have spawned end-to-end monocular VO methods in supervised [wang2017deepvo, bloesch2018codeslam, teed2019deepv2d, yang2020d3vo, wang2021tartanvo, teed2022deep, shen2023dytanvo, ye2023pvo] and unsupervised [zhou2017unsupervised, yin2018geonet, ranjan2019competitive, li2020self] settings. DeepVO [wang2017deepvo] utilizes Recurrent Neural Networks to model sequential data, while SfMLearner [zhou2017unsupervised] develops an unsupervised framework to learn the depth and camera motion from unlabeled monocular videos. TartanVO [wang2021tartanvo] proposes a VO model that generalizes to multiple datasets and real-world scenarios. DROID-SLAM [teed2021droid] follows learning-based optimization from RAFT [teed2020raft] and proposes an end-to-end system, integrating flow, confidence, and geometric optimization via iterative GRU updates. The followed-up work DPVO [teed2022deep] further cuts computational complexity by replacing dense feature tracking with sparse patch tracking.
@@
In addition to enhancing accuracy, numerous studies have focused on the detection of dynamic objects within videos. Methods based on segmentation, such as DynaSLAM [bescos2018dynaslam] and RCVD [kopf2021robust], employ semantic cues from pre-trained models to exclude dynamic regions. However, these segmentation-based methods are constrained to predefined object categories and struggle to capture real-world motion (for example, differentiating between a stationary and a moving car). Alternatively, DytanVO [shen2023dytanvo] and ParticleSfM [zhao2022particlesfm] adopt a trajectory-based motion detection approach. This method addresses dynamic scenes using estimated point track information, thereby generating reliable camera trajectories in complex motion scenarios. Nonetheless, they depend on dense optical flow estimation to capture global motion patterns. Furthermore, dynamic tracks can be detected using additional depth information for a 3D consistency check [chen2019self, zhang2022structure], or by optimization based on photo consistency [tschernezki2021neuraldiff, wu2022d, yang2023emernerf].
@@ 3.1 Preliminary: Tracking Any Point
Tracking arbitrary pixels in an image sequence is traditionally akin to estimating optical flow, where each pixel is mapped to a corresponding location in a subsequent frame. However, two-view approaches often overlook the rich temporal context within the image sequence, leading to suboptimal outcomes. Recent advancements address this by revisiting the particle video concept and improving long-term point tracking through learning-based techniques [harley2022particle, doersch2023tapir, karaev2023cotracker]. These TAP methods can track points over multiple frames and detect occlusions, significantly boosting tracking robustness under challenging scenarios.
@@
PIPs [harley2022particle], a notable learning-based TAP method, offers a promising solution featuring local cost volumes and iterative refinement. Initially, each image $\mathbf{I}_{s}\in\mathbf{I}$ is processed through a CNN feature extractor $\mathcal{F}$ to extract the feature map $\mathbf{Y_{s}}=\mathcal{F}(\mathbf{I}_{s})$. Point features $\mathbf{f}_{q}$ are computed by bilinear sampling at the query positions $\mathbf{x_{s}}$ on $\mathbf{Y}_{s}$. These point features from each image are then concatenated to form a point feature tensor $\mathbf{F}=[\mathbf{f}_{1},\ldots,\mathbf{f}_{S}]$. PIPs approaches the TAP problem by iteratively updating the state variables $(\mathbf{X},\mathbf{F})$, with the initial states obtained by duplicating the query positions and features. During the $k$-th iterative refinement, PIPs computes a multi-scale local cost volume $\mathbf{C}[\mathbf{X}^{k}]$ centered around the current estimated point position $\mathbf{X}^{k}$. Utilizing this local evidence, PIPs predicts updates for the state variables via
@@
where $\text{pos}(\cdot)$ refers to the positional embedding. The final point feature $\mathbf{F}^{K}$ is used to predict point visibility with a simple linear projection layer $\mathcal{G}_{v}$, yielding $\mathbf{V}=\mathcal{G}_{v}(\mathbf{F}^{K})$. Recent advances extend PIPs by replacing the MLP-based refiner with an attention-based refiner and adding spatial feature aggregation to track multiple queries together [karaev2023cotracker].
@@ 3.2 Anchor-based Dynamic Track Estimation
We propose an innovative method for dynamic track estimation that leverages visual, temporal, and inter-track information. While common techniques like monocular segmentation [kopf2021robust] excel at identifying moving objects using visual cues but can erroneously label static objects, like parked cars, as dynamic. Trajectory-based methods [zhao2022particlesfm], utilizing optical flows from two or multiple frames, can differentiate static and moving objects apart but tend to ignore visual information and can be inefficient due to dense trajectories computation. Merging these two distinct methods is nontrivial; however, the formulation of the TAP pipeline, enhanced with inter-track attention, offers a viable solution.
@@ 3.3 Temporal Probability Modeling
Supervised learning applied to large-scale datasets has yielded notable advancements in learning-based point tracking methods. Nonetheless, trajectory estimation errors remain a concern, especially under challenging conditions like occlusions, motion blur, or textureless regions. Recognizing the varying reliability of trajectories, integrating an uncertainty estimation mechanism into the correspondence estimation process is beneficial [chen2023uncertainty]. This is particularly crucial for quality-sensitive tasks such as visual odometry.
@@
Kernel-based Estimation. Rather than just predicting the positions of points within a trajectory, our model focuses on recovering the distribution of point trajectories by estimating its parameters. The location parameter predictions can be associated with the previous point trajectory as $\bm{\mu}_{a};\bm{\mu}_{b}]=\mathbf{X}$. However, directly deriving a symmetric and positive definite scale matrix from model outputs is challenging. To address this, we employ a kernel-based approach with the linear kernel $K(\mathbf{x},\mathbf{y})=\mathbf{x}^{T}\mathbf{y}$. During each iteration, two scale matrices $(\mathbf{\Sigma}_{a},\mathbf{\Sigma}_{b})$ are constructed by first applying two linear projection layers to the point features, represented as $\mathbf{F}_{a}^{k}=\mathcal{G}_{a}(\mathbf{F}^{k})$ and $\mathbf{F}_{b}^{k}=\mathcal{G}_{b}(\mathbf{F}^{k})$. We then compute the scale matrices as $\mathbf{\Sigma}_{a}=K(\mathbf{F}_{a},\mathbf{F}_{a})+\sigma\mathbf{I}$, and $\mathbf{\Sigma}_{b}=K(\mathbf{F}_{b},\mathbf{F}_{b})+\sigma\mathbf{I}$, where $\sigma$ is a small positive value and $\mathbf{I}$ is the identity matrix. The model parameters are refined through Maximum Likelihood Estimation (MLE) using a negative log-likelihood (NLL) loss function.
@@ 3.4 LEAP-VO
Keypoint Extraction and Tracking. In our VO system, the LEAP model serves as the front-end. When a new image frame $\mathbf{I}_{t}$ is captured at time $t$, we sample $N$ new keypoints from this image to initiate point tracking. We consistently employ the distributed image gradient-based sampling method (refer to Section 3.2) for keypoint selection due to its favorable properties in feature distinctiveness and distribution. This approach also obviates the need for additional anchors during tracking. The LEAP model then tracks these keypoints across the last $S_{LP}$ frames, performing bidirectional tracking (both forward and backward in time). Notably, $S_{LP}$ can exceed the model window $S$, thus enabling more expansive tracking capabilities through sequential prediction [harley2022particle, karaev2023cotracker].
Track Filtering. Theoretically, precise camera pose estimation between two images can be attained with just a few accurate correspondences under certain conditions. However, the quality of correspondences obtained from feature matching or optical flow can vary. Traditional VO pipelines often use RANSAC [fischler1981ransac] to filter out incorrect matches, but their efficiency decreases with more keypoints and views. Our approach, in contrast, bypasses the sampling-based method and uses trajectory quality assessments from LEAP for more efficient and informative track filtering.
@@
where $\|\cdot\|_{\rho}$ is the distance metric, and the weight $w_{i\rightarrow j,n}$ is derived from the track filtering results. We use the Gauss-Newton method [triggs1999bundle] to optimize Sec. 3.4 for $K_{BA}$ iterations. At each iteration, we compute the camera poses update $\Delta\xi^{(k)}\in\mathfrak{se}(3)~\textrm{(lie-algebra corresponding to}~\mathbf{T})$, and the depth update $\Delta\mathbf{D}^{(k)}$ for each point. The optimization can be solved efficiently with the Schur decomposition. We use the robust Huber loss function for the distance metric.
@@ 4.1 Datasets and Metrics
Replica. Replica [straub2019replica] provides synthetic indoor environments for simulating the image-capturing process with a moving camera. We use the camera trajectories provided by [zhi2021place] for the localization task, resulting in 16 camera trajectories derived from 8 static scenes. Each sequence comprises 900 frames with rendered RGB images. We use “Sequence 1” for the visual odometry evaluation.
MPI Sintel. The MPI Sintel [Butler2012sintel] dataset contains dynamic image sequences from open-source 3D animated short films. These sequences consist of large and complex object motion, along with other imaging effects such as motion blur and defocus blur. Following prior works [kopf2021robust, zhao2022particlesfm], we assess dynamic VO performance on the MPI Sintel dataset, with each sequence containing 20 to 50 image frames.
TartanAir Shibuya. The TartanAir Shibuya dataset [qiu2022airdos] features dynamic scenes from two scenarios: “Standing Humans” and “Road Crossing”. The majority of the humans remain stationary in ”Standing Humans” sequences, while ”Road Crossing” sequences provide a more challenging setting where multiple humans are moving in different directions. Each sequence in the dataset includes 100 frames with more than 30 tracked moving humans.
Metrics. For visual odometry, we compare the Absolute Translation Error (ATE), Relative Translation Error (RPE trans), and Relative Rotation Error (RPE rot) following  [zhao2022particlesfm, zhang2022structure, teed2022deep]. ATE measures how much the estimated trajectory deviated from the ground truth trajectory, which is calculated as the root mean square error (RMSE) overall all pose transformations after scaling and alignment. “RPE trans” computes translation errors made over a certain distance (meter), and “RPE rot” computes the estimated rotation over a certain distance (degree). Both “RPE trans” and “RPE rot” are calculated on all poses and then averaged.
@@
ORB-SLAM2 [murorbslam2] 0.086 0.030 0.650
DynaSLAM [bescos2018dynaslam] 0.039 0.017 0.366
DROID-SLAM [teed2021droid] 0.267 0.036 2.631
TartanVO [wang2021tartanvo] 0.406 0.036 2.063
DytanVO [shen2023dytanvo] 0.289 0.035 2.146
DPVO [teed2022deep] 0.257 0.036 2.635
@@
Table 1: Camera tracking results on Replica [straub2019replica]. The statistics demonstrate that our method achieves better results than other competitive baselines.
@@
ORB-SLAM2 [murorbslam2] X X X
DynaSLAM [bescos2018dynaslam] X X X
DROID-SLAM [teed2021droid] 0.175 0.084 1.912
TartanVO [wang2021tartanvo] 0.238 0.093 1.305
DytanVO [shen2023dytanvo] 0.131 0.097 1.538
DPVO [teed2022deep] 0.076 0.078 1.722
@@
Table 2: Camera tracking results on MPI Sintel [Butler2012sintel]. The statistics show that our method outperforms other competitive baselines.‘X’ denotes the tracking failure case.
@@
ORB-SLAM2 [murorbslam2] w/ mask 0.0788 0.0060 0.0657 0.0196 0.0148 1.0984 0.8476 0.3044
DynaSLAM [bescos2018dynaslam] X 0.8836 0.3907 0.4196 0.4925 0.6446 0.6539 (0.5808)
AirDOS [qiu2022airdos] w/ mask 0.0606 0.0193 0.0951 0.0331 0.0206 0.2230 0.5625 0.1449
DROID-SLAM [teed2021droid] 0.0051 0.0073 0.0103 0.0120 0.2278 0.0253 0.5788 0.1238
TartanVO [wang2021tartanvo] 0.0600 0.1605 0.2762 0.1814 0.2174 0.3228 0.5009 0.2456
DytanVO [shen2023dytanvo] 0.0327 0.1017 0.0608 0.0516 0.0755 0.0365 0.0660 0.0607
DPVO [wang2017deepvo] 0.0539 0.1603 0.1788 0.2858 0.1196 0.1119 0.1498 0.1514
@@
Table 3: ATE (m) results on TartanAir-Shibuya Sequences [qiu2022airdos]. $(\cdot)$ denotes averaging on valid scenes only. We can observe that our method shows better results for ATE on most sequences, outperforming all SLAM and VO baselines on average ATE.
@@
Figure 5: Qualitative results of camera trajectory estimation on TartanAir-Shibuya [qiu2022airdos]. The visualizations show that our method provides more robust and accurate camera pose trajectories, especially in hard cases (RoadCrossing 06 and RoadCrossing 07).